# Unlearning Inversion Attacks for Graph Neural Networks

### v1

@@ 7 Conclusion
  • Section A: Additional Related Works.
  • Section B: Model Details.
  • Section C: Theoretical Motivation of the Confidence Pitfall.
  • Section D: Missing Proofs in Section C.
  • Section E: Preliminary Experiments.
@@
  • Section H: Impact Statement.
Appendix A Additional Related Works
@@ A.1 Machine Unlearning
Machine unlearning aims to remove the influence of specific training samples from a trained model, balancing utility, removal guarantees, and efficiency bourtoule2021machine ; liu2024breaking . Unlike full retraining, unlearning seeks practical alternatives to efficiently revoke data. These methods can help removing data effect in simple learning settings like statistical query learning cao2015towards , and recently have broad applications in LLMs yao2024large ; liu2024large ; liu2025rethinking , generative models li2024machine ; zhang2024unlearncanvas ; zhang2024defensive , e-commerce chen2022recommendation ; li2023ultrare ; zhang2024recommendation , and graph learning chen2022graph ; wu2023gif ; dong2024idea .
@@
Existing machine unlearning methods broadly fall into two categories: exact and approximate unlearning. Exact unlearning methods, such as SISA bourtoule2021machine , partition data into shards, train sub-models in each shard independently, and merge them together, which allows targeted retraining for removing certain data from a shard. ARCANE yan2022arcane improves this by framing unlearning as one-class classification and caching intermediate states, enhancing retraining efficiency. These ideas have been extended to other domains, including graph learning chen2022graph ; wang2023inductive and ensemble methods brophy2021machine . However, exact methods often degrade performance due to weak sub-models and straightforward ensembling.
@@
Approximate unlearning methods update model parameters to emulate removal effects with better efficiency-performance trade-offs. For instance, Jia et al. jia2023model uses pruning and sparsity-regularized fine-tuning to approximate exact removal. Tarun et al. tarun2023fast propose a model-agnostic class removal technique using error-maximizing noise and a repair phase. Liu et al. liu2023muter tackle adversarially trained models with a closed-form Hessian-guided update, approximated efficiently without explicit inversion.
A.2 Privacy Vulnerabilities of Machine Unlearning
@@
A growing line of research investigates the privacy vulnerabilities of machine unlearning. Chen et al. chen2021machine show that if an adversary has access to both the pre- and post-unlearning black-box model outputs, they can infer the membership status of a target sample. A recent work thudi2022necessity highlights the limitations of approximate unlearning by arguing that unlearning is only well-defined at the algorithmic level, and parameter-level manipulations alone may be insufficient. Unlearning inversion attacks hu2024learn go further by reconstructing both features and labels of training samples using two model versions, while Bertran et al. bertran2024reconstruction provide an in-depth analysis of such reconstruction attacks in regression settings.
@@
These findings motivate our investigation into unlearned GNNs, where we aim to capture residual signals left by the unlearned data. While prior attacks mostly target general-purpose models and focus on red-teaming under strong assumptions, such as access to both pre- and post-unlearning models chen2021machine ; hu2024learn ; bertran2024reconstruction , our work considers a more practical black-box setting motivated by real-world social network scenarios. Furthermore, our attack explicitly accounts for the unique structural and relational properties of graphs, in contrast to prior work primarily focused on i.i.d. data.
A.3 Graph Unlearning
@@
Graph Unlearning enables the efficient removal of unwanted data’s influence from trained graph ML models chen2022graph ; said2023survey ; fan2025opengu . This removal process balances model utility, unlearning efficiency, and removal guarantees, following two major lines of research: retrain-based unlearning and approximate unlearning.
@@
Retrain-based unlearning partitions the original training graph into disjoint subgraphs, training independent submodels on them, enabling unlearning through retraining on a smaller subset of the data. Specifically, GraphEraser chen2022graph pioneered the first retraining-based unlearning framework for GNNs, utilizing balanced clustering methods for subgraph partitioning and ensembling submodels in prediction with trainable fusion weights to enhance model utility. GUIDE wang2023inductive presents an important follow-up by extending the retraining-based unlearning paradigm to inductive graph learning settings. Subsequently, many studies wang2023inductive ; li2023ultrare ; zhang2024graph ; li2025community have made significant contributions to improving the Pareto front of utility and efficiency in these methods, employing techniques such as data condensation li2025community and enhanced clustering li2023ultrare ; zhang2024graph .
@@
Approximate unlearning efficiently updates model parameters to remove unwanted data. Certified graph unlearning chien2022certified provides an important early exploration of approximate unlearning in SGC wu2019simplifying , with provable unlearning guarantees. GraphGuard wu2024graphguard introduces a comprehensive system to mitigate training data misuse in GNNs, featuring a significant gradient ascent unlearning method as one of its core components. These gradient ascent optimization objectives can be approximated using Taylor expansions, inspiring several beautiful influence function-based methods koh2017understanding ; wu2023gif ; wu2023certified . Specifically, GIF wu2023gif presents a novel influence function-based unlearning approach tailored to graph data, considering feature, node, and edge unlearning settings, while CEU wu2023certified focuses on edge unlearning and establishes theoretical bounds for removal guarantees. Building on this, IDEA dong2024idea proposes a theoretically sound and flexible approximate unlearning framework, offering removal guarantees across node, edge, and feature unlearning settings. Recent innovative works have further advanced the scalability pan2023unlearning ; li2024tcgu ; yi2025scalable ; yang2025erase ; zhang2025dynamic and model utility li2024towards ; zhang2025node of approximate unlearning methods, pushing the boundaries of promising applications. Additionally, a noteworthy contribution is GNNDelete cheng2023gnndelete , which unlearns graph knowledge by employing an intermediate node embedding mask between GNN layers. This method offers a unique solution that differs from both retraining and model parameter modification approaches.
@@
A.4 Membership Inference Attack for GNNs
Membership Inference Attack (MIA) is a privacy attack targeting ML models, aiming to distinguish whether a specific data point belongs to the training set shokri2017membership ; hu2022membership ; dai2023unified . Such attacks have profoundly influenced ML across various downstream applications, including computer vision chen2020gan ; choquette2021label , NLP mattern2023membership ; meeus2024did , and recommender systems zhang2021membership ; yuan2023interaction . Recently, MIA has been extended to graph learning, where a pioneering work duddu2020quantifying explored the feasibility of membership inference in classical graph embedding models. Subsequently, interest has shifted towards attacking graph neural networks (GNNs), with several impactful and innovative studies revealing GNNs’ privacy vulnerabilities in node classification he2021node ; he2021stealing ; olatunji2021membership ; zhang2022inference and graph classification tasks wu2021adapting ; lin2024stealing , covering cover node-level he2021node ; olatunji2021membership , link-level he2021stealing , and graph-level wu2021adapting ; zhang2022inference inference risks. Building on this, GroupAttack zhang2023demystifying presents a compelling advancement in link-stealing attacks he2021stealing on GNNs, theoretically demonstrating that different edge groups exhibit varying risk levels and require distinct attack thresholds, while a label-only attack has been proposed to target node-level privacy vulnerabilities conti2022label with a stricter setting. Another significant line of research involves graph model inversion attacks, which aim to reconstruct the graph structure using model gradients from white-box models zhang2021graphmi or approximated gradients from black-box models zhang2022model .
@@ A.4 Membership Inference Attack for GNNs
Appendix B Model Details
@@
B.1 Shadow Training Algorithm for TrendAttack
@@
4:Train the shadow victim model on $\mathcal{G}^{\mathrm{sha}}_{\mathrm{orig}}$ with Eq. (1) to obtain $\boldsymbol{\theta}_{\mathrm{orig}}^{\prime}$triangleright Train victim model
@@
6:Unlearn $\Delta\mathcal{E}^{\mathrm{sha}}$ with Eq. (2) to obtain $\boldsymbol{\theta}_{\mathrm{un}}^{\prime}$triangleright Unlearn victim model
@@
9:Randomly select some existing edges from $\mathcal{E}^{\mathrm{sha}}_{\mathrm{un}}$ to obtain $\mathcal{Q}^{+}_{\mathrm{mem}}$triangleright $\mathcal{Q}^{+}_{\mathrm{mem}}\subseteq\mathcal{E}^{\mathrm{sha}}_{\mathrm{un}}$
10:Randomly select some negative edges $\mathcal{Q}^{-}$triangleright $\mathcal{Q}^{-}\cap\mathcal{E}^{\mathrm{sha}}_{\mathrm{un}}=\emptyset$
@@
B.2 The Attack Process of TrendAttack
@@
1:Unlearned graph $\mathcal{G}_{\mathrm{un}}$, black-box unlearned model $f_{\mathcal{G}_{\mathrm{un}}}(\cdot;\boldsymbol{\theta}_{\mathrm{un}})$, query set $\mathcal{Q}$, trained attack model $g(\cdot,\cdot)$, trend feature order $k$
@@
5:$\mathcal{P}_{\mathcal{Q}}\leftarrow\emptyset,\mathcal{F}_{\mathcal{Q}}\leftarrow\emptyset$
@@
7:     $\mathcal{F}_{\mathcal{Q}}\leftarrow\mathcal{F}_{\mathcal{Q}}\cup\{(v_{i},\mathbf{x}_{i})\}$triangleright Feature knowledge
@@
10:         $\mathbf{p}_{j}\leftarrow f_{\mathcal{G}_{\mathrm{un}}}(v_{j};\boldsymbol{\theta}_{\mathrm{un}})$
@@
16:     Compute trend features $\widetilde{\boldsymbol{\tau}}_{i}$ for node $v_{i}$ with Eq. (3) and Eq. (4)
@@
In this algorithm, we first request the partial knowledge $\mathcal{K}_{\mathcal{Q}}$ from the unlearned graph $\mathcal{G}_{\mathrm{un}}$ and the black-box victim model $f_{\mathcal{G}_{\mathrm{un}}}(\cdot;\boldsymbol{\theta}_{\mathrm{un}})$ for our links of interest $\mathcal{Q}$ (lines 1–11). This process is highly flexible and can effortlessly incorporate many different levels of the attacker’s knowledge. For instance, the feature knowledge $\mathcal{F}_{\mathcal{Q}}$ (line 5) is optional if the model API owner does not respond to node feature requests on the unlearned graph $\mathcal{G}_{\mathrm{un}}$.
@@
Appendix C Theoretical Motivation of the Confidence Pitfall
In this section, we present our main theoretical results, offering a robust theoretical foundation for our model design, with a specific focus on the confidence pitfall intuition in Section 5.1.
Our analysis of the probability similarity gap (Claim 5.1) suggests that unlearned edges $\mathcal{Q}^{+}_{\mathrm{un}}$ and remaining membership edges $\mathcal{Q}^{+}_{\mathrm{mem}}$ require different probability-similarity thresholds for accurate inference. The challenge, however, is that the attacker cannot directly tell whether a given query $(v_{i},v_{j})$ belongs to $\mathcal{Q}^{+}_{\mathrm{un}}$ or $\mathcal{Q}^{+}_{\mathrm{mem}}$.
To address this, we analyze how removing a specific edge $(v_{i},v_{j})$ affects model outputs, using the widely used analytical framework of influence functions koh2017understanding ; wu2023gif ; wu2023certified . This analysis separates two effects: (i) the immediate impact of dropping the edge from the training graph $\mathcal{G}_{\mathrm{orig}}$, and (ii) the subsequent adjustment of model parameters by the unlearning procedure. We focus on a linear GCN model, which, despite its simplicity, captures the core behavior of many GNN architectures (see Remark D.2 in Appendix D).
Specifically, we first present the influence result of all the unlearned edges $\Delta\mathcal{E}$ as follows:
Theorem C.1 (Closed-form Edge Influence, Informal).
Let $f(\mathbf{C},\mathbf{X};\mathbf{w}^{\star})$ be a linear GCN with propagation matrix $\mathbf{C}\in\mathbb{R}^{n\times n}$ and parameters $\mathbf{w}^{\star}$ obtained by least-squares on labels $\mathbf{y}\in\mathbb{R}^{d}$. If we denote the matrix of all node predictions by $\mathbf{P}$ and let $\boldsymbol{\Xi}:=\sum_{(v_{i},v_{j})\in\Delta\mathcal{E}}\mathbf{e}_{i}\mathbf{e}_{j}^{\top}$ be the adjacency matrix of the removed edges, then unlearning $\Delta\mathcal{E}$ produces:
  • Weight influence:
    $\mathcal{I}(\mathbf{w}^{\star})=\mathbf{H}^{-1}(\mathbf{X}^{\top}\boldsymbol{\Xi}^{\top}\mathbf{y}-\mathbf{X}^{\top}\boldsymbol{\Xi}^{\top}\mathbf{C}\mathbf{X}\mathbf{w}^{\star}-\mathbf{X}^{\top}\mathbf{C}^{\top}\boldsymbol{\Xi}\mathbf{X}\mathbf{w}^{\star}),$
  • Model output influence:
    $\mathcal{I}(\mathbf{P})=\boldsymbol{\Xi}\mathbf{X}\mathbf{w}^{\star}+\mathbf{C}\mathbf{X}\mathbf{H}^{-1}\Bigl{(}\mathbf{X}^{\top}\boldsymbol{\Xi}^{\top}\mathbf{y}-\mathbf{X}^{\top}\mathbf{C}^{\top}\boldsymbol{\Xi}\mathbf{X}\mathbf{w}^{\star}-\mathbf{X}^{\top}\boldsymbol{\Xi}^{\top}\mathbf{C}\mathbf{X}\mathbf{w}^{\star}\Bigr{)},$
where $\mathbf{H}$ is the Hessian of the least squares loss evaluated at $\mathbf{w}^{\star}$.
Proof.
This follows directly from Theorem D.8 and Theorem D.9. See Appendix D for proofs. ∎
While Theorem C.1 characterizes the combined effect of all unlearned edges, we also need to understand the effect of a single edge on a single node’s output. This finer-grained analysis reveals why nodes adjacent to unlearned edges exhibit a distinctive drop in confidence compared to more distant nodes.
Corollary C.2 (Single-Edge to Single-Node Influence, Informal).
Under the same setup as Theorem C.1, the influence of an undirected edge $(v_{i},v_{j})\in\Delta\mathcal{E}$ on the output $\mathbf{p}_{k}$ of node $v_{k}\in\mathcal{V}$ can be decomposed as follows:
$\small\mathcal{I}(\mathbf{p}_{k})=$ $~{}\underbrace{\boldsymbol{1}{\{v_{k}=v_{i}\}}\cdot(\mathbf{x}^{\top}_{i}\mathbf{w}^{\star})+\boldsymbol{1}{\{v_{k}=v_{j}\}}\cdot(\mathbf{x}^{\top}_{j}\mathbf{w}^{\star})}_{\mathrm{edge~{}influence}}$
$~{}~{}~{}~{}-\underbrace{\langle(\mathbf{x}_{j}^{\top}\mathbf{w}^{\star})\mathbf{z}_{i}+(\mathbf{x}_{i}^{\top}\mathbf{w}^{\star})\mathbf{z}_{j},\mathbf{z}_{k}\rangle_{\mathbf{H}^{-1}}}_{\mathrm{magnitude~{}weight~{}influence}}+\underbrace{\langle(y_{j}-\mathbf{z}_{j}^{\top}\mathbf{w}^{\star})\mathbf{x}_{i}+(y_{i}-\mathbf{z}_{i}^{\top}\mathbf{w}^{\star})\mathbf{x}_{j},\mathbf{z}_{k}\rangle_{\mathbf{H}^{-1}}}_{\mathrm{error~{}weight~{}influence}},$
where $\mathbf{Z}:=\mathbf{C}\mathbf{X}$ and $\mathbf{z}_{l}$ is the $l$-th row of $\mathbf{Z}$, and $\boldsymbol{1}{\{\cdot\}}$ denotes the indicator function.
Proof.
Please see Corollary D.11 in Appendix D. ∎
In the corollary above, the third error weight influence term is governed by the empirical residuals at the optimal weight, $(y_{i}-\mathbf{z}_{i}^{\top}\mathbf{w}^{\star})$ and $(y_{j}-\mathbf{z}_{j}^{\top}\mathbf{w}^{\star})$. Under the assumption of a well-trained (i.e., learnable) graph ML problem, these residuals become negligibly small, making the error weight influence term nearly zero.
Meanwhile, the first edge influence term activates only when $v_{k}$ coincides with one of the endpoints $v_{i}$ or $v_{j}$, producing a direct and large perturbation. The second magnitude weight influence term, which depends on the inner-product similarity $\langle(\mathbf{x}_{j}^{\top}\mathbf{w}^{\star})\mathbf{z}_{i}+(\mathbf{x}_{i}^{\top}\mathbf{w}^{\star})\mathbf{z}_{j},\mathbf{z}_{k}\rangle_{\mathbf{H}^{-1}}$, is also relatively larger when $v_{k}\in\{v_{i},v_{j}\}$ and smaller otherwise. As a result, for $v_{k}=v_{i}$ or $v_{j}$, the sum of a substantial positive edge influence and a significant negative magnitude weight influence yields a dramatic net effect on the model output. Therefore, the influence on the endpoints of unlearned edges may be more significant than on other nodes. Based on this, we reasonably assume their confidence will drop and make the following claim:
Claim C.3 (Confidence Pitfall, Restatement of Claim 5.2).
$\mathrm{AvgConf}(\mathcal{V}_{\mathcal{Q}^{+}_{\mathrm{un}}})<\mathrm{AvgConf}(\mathcal{V}\setminus\mathcal{V}_{\mathcal{Q}^{+}_{\mathrm{un}}}),$
where $\mathcal{V}_{\mathcal{Q}_{0}}:=\{v_{i}:(v_{i},v_{j})\in\mathcal{Q}_{0}~{}\mathrm{or}~{}(v_{j},v_{i})\in\mathcal{Q}_{0}\}$ is the set of all nodes involved in query subset $\mathcal{Q}_{0}$, and $\mathrm{AvgConf}(\mathcal{V}_{0}):=\frac{1}{|\mathcal{V}_{0}|}\sum_{v_{i}\in\mathcal{V}_{0}}\max_{\ell}\mathbf{p}_{i,\ell}$ is the average model confidence within a node set $\mathcal{V}_{0}$.
Further empirical evidence supporting this claim can be found in Appendix E.
Appendix D Missing Proofs in Section C
@@
In this section, we provide formal definitions for all concepts introduced in Section C and supplement the missing technical proofs. We begin by describing the architecture and training process of linear GCN, and then compute the edge influence on model weights. Next, we analyze how edge influence affects the model output, considering both the direct impact from the edge itself and the indirect influence mediated through model weights.
D.1 Linear GCN

### v2

@@ Acknowledgments
  • Section A: Notations.
  • Section B: Additional Related Works.
  • Section C: Model Details.
  • Section D: Missing Proofs in Section 5.1.
  • Section E: Preliminary Study.
@@
  • Section H: Discussions.
Appendix A Notations
In this paper, calligraphy uppercase letters (e.g., $\mathcal{X}$) denote sets, bold uppercase letters (e.g., $\mathbf{X}$) denote matrices, bold lowercase letters (e.g., $\mathbf{x}$) denote column vectors, and normal letters (e.g., $x$) indicate scalars. Let $\mathbf{e}_{u}\in\mathbb{R}^{d}$ be the column vector with the $u$-th element as $1$ and all others as 0. We use $||$ to denote concatenating two vectors. We define the weighted inner product with a PSD matrix $\mathbf{H}$ as $\langle\mathbf{x},\mathbf{y}\rangle_{\mathbf{H}}:=\mathbf{x}^{\top}\mathbf{H}\mathbf{y}$. We use $\bm{1}\{\cdot\}$ to denote the indicator function, which returns 1 if the condition in brackets is satisfied and 0 otherwise. We use $\mathcal{A}\setminus\mathcal{B}:=\{x:x\in\mathcal{A},x\notin\mathcal{B}\}$ to denote the set difference between sets $\mathcal{A}$ and $\mathcal{B}$.
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ denote a graph, where $\mathcal{V}=\{v_{1},\cdots,v_{n}\}$ is the node set and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the edge set. The node feature of node $v_{i}\in\mathcal{V}$ is denoted by $\mathbf{x}_{i}\in\mathbb{R}^{d}$, and the feature matrix for all nodes is denoted by $\mathbf{X}=[\mathbf{x}_{1},\cdots,\mathbf{x}_{n}]^{\top}\in\mathbb{R}^{n\times d}$. The adjacency matrix $\mathbf{A}\in\{0,1\}^{n\times n}$ encodes the edge set, where $A_{i,j}=1$ if $(v_{i},v_{j})\in\mathcal{E}$, and $\mathbf{A}_{i,j}=0$ otherwise. Specifically, $\mathbf{D}\in\mathbb{R}^{n\times n}$ is the degree matrix, where the diagonal elements $D_{i,i}=\sum_{j=1}^{n}A_{i,j}$. We use $\mathcal{N}(v)$ to denote the neighborhood of node $v_{i}\in\mathcal{V}$, and use $\widehat{\mathcal{N}}(v)$ to denote a subset of $\mathcal{N}(v)$. Specifically, $\mathcal{N}^{(k)}(v_{i})$ represents all nodes in $v_{i}$’s $k$-hop neighborhood, where $\mathcal{N}^{(0)}(v_{i}):=\{v_{i}\}$ and for $k\geq 1$, $\mathcal{N}^{(k)}(v_{i}):=\{v_{j}:(v_{i},v_{j})\in\mathcal{E},v_{i}\in\mathcal{N}^{(k-1)}(v_{i})\}\cup\mathcal{N}^{(k-1)}(v_{i})$.
Appendix B Additional Related Works
@@ B.1. Machine Unlearning
Machine unlearning aims to remove the influence of specific training samples from a trained model, balancing utility, removal guarantees, and efficiency (bourtoule2021machine; liu2024breaking). Unlike full retraining, unlearning seeks practical alternatives to efficiently revoke data. These methods can help removing data effect in simple learning settings like statistical query learning (cao2015towards), and recently have broad applications in LLMs (yao2024large; liu2024large; liu2025rethinking), generative models (li2024machine; zhang2024unlearncanvas; zhang2024defensive), e-commerce (chen2022recommendation; li2023ultrare; zhang2024recommendation), and graph learning (chen2022graph; wu2023gif; dong2024idea).
@@
Existing machine unlearning methods broadly fall into two categories: exact and approximate unlearning. Exact unlearning methods, such as SISA (bourtoule2021machine), partition data into shards, train sub-models in each shard independently, and merge them together, which allows targeted retraining for removing certain data from a shard. ARCANE (yan2022arcane) improves this by framing unlearning as one-class classification and caching intermediate states, enhancing retraining efficiency. These ideas have been extended to other domains, including graph learning (chen2022graph; wang2023inductive) and ensemble methods (brophy2021machine). However, exact methods often degrade performance due to weak sub-models and straightforward ensembling.
@@
Approximate unlearning methods update model parameters to emulate removal effects with better efficiency-performance trade-offs. For instance, Jia et al. (jia2023model) uses pruning and sparsity-regularized fine-tuning to approximate exact removal. Tarun et al. (tarun2023fast) propose a model-agnostic class removal technique using error-maximizing noise and a repair phase. Liu et al. (liu2023muter) tackle adversarially trained models with a closed-form Hessian-guided update, approximated efficiently without explicit inversion.
B.2. Privacy Vulnerabilities of Machine Unlearning
@@
A growing line of research investigates the privacy vulnerabilities of machine unlearning. Chen et al. (chen2021machine) show that if an adversary has access to both the pre- and post-unlearning black-box model outputs, they can infer the membership status of a target sample. A recent work (thudi2022necessity) highlights the limitations of approximate unlearning by arguing that unlearning is only well-defined at the algorithmic level, and parameter-level manipulations alone may be insufficient. Unlearning inversion attacks (hu2024learn) go further by reconstructing both features and labels of training samples using two model versions, while Bertran et al. (bertran2024reconstruction) provide an in-depth analysis of such reconstruction attacks in regression settings.
@@
These findings motivate our investigation into unlearned GNNs, where we aim to capture residual signals left by the unlearned data. While prior attacks mostly target general-purpose models and focus on red-teaming under strong assumptions, such as access to both pre- and post-unlearning models (chen2021machine; hu2024learn; bertran2024reconstruction), our work considers a more practical black-box setting motivated by real-world social network scenarios. Furthermore, our attack explicitly accounts for the unique structural and relational properties of graphs, in contrast to prior work primarily focused on i.i.d. data.
B.3. Graph Unlearning
@@
Graph Unlearning enables the efficient removal of unwanted data’s influence from trained graph ML models (chen2022graph; said2023survey; fan2025opengu). This removal process balances model utility, unlearning efficiency, and removal guarantees, following two major lines of research: retrain-based unlearning and approximate unlearning.
@@
Retrain-based unlearning partitions the original training graph into disjoint subgraphs, training independent submodels on them, enabling unlearning through retraining on a smaller subset of the data. Specifically, GraphEraser (chen2022graph) pioneered the first retraining-based unlearning framework for GNNs, utilizing balanced clustering methods for subgraph partitioning and ensembling submodels in prediction with trainable fusion weights to enhance model utility. GUIDE (wang2023inductive) presents an important follow-up by extending the retraining-based unlearning paradigm to inductive graph learning settings. Subsequently, many studies (wang2023inductive; li2023ultrare; zhang2024graph; li2025community) have made significant contributions to improving the Pareto front of utility and efficiency in these methods, employing techniques such as data condensation (li2025community) and enhanced clustering (li2023ultrare; zhang2024graph).
@@
Approximate unlearning efficiently updates model parameters to remove unwanted data. Certified graph unlearning (chien2022certified) provides an important early exploration of approximate unlearning in SGC (wu2019simplifying), with provable unlearning guarantees. GraphGuard (wu2024graphguard) introduces a comprehensive system to mitigate training data misuse in GNNs, featuring a significant gradient ascent unlearning method as one of its core components. These gradient ascent optimization objectives can be approximated using Taylor expansions, inspiring several beautiful influence function-based methods (koh2017understanding; wu2023gif; wu2023certified). Specifically, GIF (wu2023gif) presents a novel influence function-based unlearning approach tailored to graph data, considering feature, node, and edge unlearning settings, while CEU (wu2023certified) focuses on edge unlearning and establishes theoretical bounds for removal guarantees. Building on this, IDEA (dong2024idea) proposes a theoretically sound and flexible approximate unlearning framework, offering removal guarantees across node, edge, and feature unlearning settings. Recent innovative works have further advanced the scalability (pan2023unlearning; li2024tcgu; yi2025scalable; yang2025erase), model utility (li2024towards; zhang2025node), and dynamic adaptivity (zhang2025dynamic) of approximate unlearning methods, pushing the boundaries of promising applications. Additionally, a noteworthy contribution is GNNDelete (cheng2023gnndelete), which unlearns graph knowledge by employing an intermediate node embedding mask between GNN layers. This method offers a unique solution that differs from both retraining and model parameter modification approaches.
@@
B.4. Membership Inference Attack for GNNs
Membership Inference Attack (MIA) is a privacy attack targeting ML models, aiming to distinguish whether a specific data point belongs to the training set (shokri2017membership; hu2022membership; dai2023unified; liu2025exposing). Such attacks have profoundly influenced ML across various downstream applications, including computer vision (chen2020gan; choquette2021label), NLP (mattern2023membership; meeus2024did), and recommender systems (zhang2021membership; yuan2023interaction). Recently, MIA has been extended to graph learning, where a pioneering work (duddu2020quantifying) explored the feasibility of membership inference in classical graph embedding models. Subsequently, interest has shifted towards attacking graph neural networks (GNNs), with several impactful and innovative studies revealing GNNs’ privacy vulnerabilities in node classification (he2021node; he2021stealing; olatunji2021membership; zhang2022inference) and graph classification tasks (wu2021adapting; lin2024stealing), covering cover node-level (he2021node; olatunji2021membership), link-level (he2021stealing), and graph-level (wu2021adapting; zhang2022inference) inference risks. Building on this, GroupAttack (zhang2023demystifying) presents a compelling advancement in link-stealing attacks (he2021stealing) on GNNs, theoretically demonstrating that different edge groups exhibit varying risk levels and require distinct attack thresholds, while a label-only attack has been proposed to target node-level privacy vulnerabilities (conti2022label) with a stricter setting. Another significant line of research involves graph model inversion attacks, which aim to reconstruct the graph structure using model gradients from white-box models (zhang2021graphmi) or approximated gradients from black-box models (zhang2022model).
@@ B.4. Membership Inference Attack for GNNs
Appendix C Model Details
@@
C.1. Shadow Training Algorithm for TrendAttack
@@
4:Train the shadow victim model on $\mathcal{G}^{\mathrm{sha}}_{\mathrm{orig}}$ with Eq. (1) to obtain $\bm{\theta}_{\mathrm{orig}}^{\prime}$ $\triangleright$ Train victim model
@@
6:Unlearn $\Delta\mathcal{E}^{\mathrm{sha}}$ with Eq. (3) to obtain $\bm{\theta}_{\mathrm{un}}^{\prime}$ $\triangleright$ Unlearn victim model
@@
9:Randomly select some existing edges from $\mathcal{E}^{\mathrm{sha}}_{\mathrm{un}}$ to obtain $\mathcal{Q}^{+}_{\mathrm{mem}} \triangleright$ $\mathcal{Q}^{+}_{\mathrm{mem}}\subseteq\mathcal{E}^{\mathrm{sha}}_{\mathrm{un}}$
10:Randomly select some negative edges $\mathcal{Q}^{-}$ $\triangleright$ $\mathcal{Q}^{-}\cap\mathcal{E}^{\mathrm{sha}}_{\mathrm{un}}=\emptyset$
@@
C.2. The Attack Process of TrendAttack
@@
1:Unlearned graph $\mathcal{G}_{\mathrm{un}}$, black-box unlearned model $f_{\mathcal{G}_{\mathrm{un}}}(\cdot;\bm{\theta}_{\mathrm{un}})$, query set $\mathcal{Q}$, trained attack model $g(\cdot,\cdot)$, trend feature order $k$
@@
[…]
@@
7:  $\mathcal{F}_{\mathcal{Q}}\leftarrow\mathcal{F}_{\mathcal{Q}}\cup\{(v_{i},\mathbf{x}_{i})\}$ $\triangleright$ Feature knowledge
@@
10:   $\mathbf{p}_{j}\leftarrow f_{\mathcal{G}_{\mathrm{un}}}(v_{j};\bm{\theta}_{\mathrm{un}})$
@@
16:  Compute trend features $\widetilde{\bm{\tau}}_{i}$ for node $v_{i}$ with Eq. (5.2.1) and Eq. (4)
@@
In this algorithm, we first request the partial knowledge $\mathcal{K}_{\mathcal{Q}}$ from the unlearned graph $\mathcal{G}_{\mathrm{un}}$ and the black-box victim model $f_{\mathcal{G}_{\mathrm{un}}}(\cdot;\bm{\theta}_{\mathrm{un}})$ for our links of interest $\mathcal{Q}$ (lines 1–11). This process is highly flexible and can effortlessly incorporate many different levels of the attacker’s knowledge. For instance, the feature knowledge $\mathcal{F}_{\mathcal{Q}}$ (line 5) is optional if the model API owner does not respond to node feature requests on the unlearned graph $\mathcal{G}_{\mathrm{un}}$.
@@
Appendix D Missing Proofs in Section 5.1
@@
In this section, we provide formal definitions for all concepts introduced in Section 5.1 and supplement the missing technical proofs. We begin by describing the architecture and training process of linear GCN, and then compute the edge influence on model weights. Next, we analyze how edge influence affects the model output, considering both the direct impact from the edge itself and the indirect influence mediated through model weights.
D.1. Linear GCN